1,2. On Yahoo! Finance, look up two stock’s betas.

McDonald’s:

4.9

Apple:

1.29

3.4. If the expected market return is 10% and the risk-free rate is 1%, what are the expected returns of the two stocks in 1 and 2 above?

5. If the expected market return is 11% and the risk-free rate is 2.50%, what is the expected return of a stock with a beta of -0.1?
Answer #1

Using CAMP model,

Expected return of McDonalds = 0.01 + 4.9(0.10 - 0.01)

Expected return of McDonalds = 45.10%

Expected return of Apple = 0.01 + 1.29(0.10 - 0.01)

Expected return of Apple = 12.61%

Expected return of stock with beta -0.1 = 0.02 -0.10(0.11 - 0.025)

Expected return = 1.65%